A current is going 3 mp, boat goes 6 miles upstream in the same amount of time it takes a boat to go 10 miles downstream. Find the speed of the boat in still water.
I dont know of which (if any) formula to use. Distance formula?
i know that the first boat only travels 6 miles in the same time because it is going against the current
i dont know if i should put the boat having a motor into concideration because if no then it took the boat 2 hours going 3 mph to go 6 miles? <-- i dont think that parts right...but..and 3.3hr to go 10 miles down stream.

Answer by texttutoring(324)

Yeah, use Distance formula D=VT, but isolate it for T: T=D/V

Because the times are the same, you can set the upstream time and downstream time equal to each other. Using Dup and Vup for upstream, and Ddown and Vdown for downstream, we know that Vup = V - 3, and Vdown = V + 3 (you subtract the current when you go against it, and add the current when it helps you):

Dup/Vup = Ddown / Vdown
6/V-3 = 10/V+3

Cross multiply and solve for V, the boat's still water velocity:
6(V+3) = 10(V-3)
6V + 18 = 10V - 30
48 = 4V
V = 12

The boat's speed in still water is 12mph.